OpenStudy (anonymous):

hi, who knows the proof of 0 * X=0

OpenStudy (anonymous):

This is the zero identity of multiplication over whatever field X belongs to.

OpenStudy (anonymous):

zero property***

OpenStudy (anonymous):

I need a proof to this simple expression

OpenStudy (sirm3d):

0x = (-1 + 1)x = (-1)x + 1(x) = -x + x = 0

OpenStudy (anonymous):

thanks sirm3d that what I exactly need

OpenStudy (sirm3d):

yw